Set up one example question (also provide solutions) which falls under the following.
Level 1:mathematical knowledge.
Level 2:routine procedures (in a real life context)
Level 3:complex procedures(in real life context)
Level 4:problem solving(in real life context?)
1
Expert's answer
2020-07-03T16:28:12-0400

Level 1. Mathematical knowledge

Question . Which of the following fractions is the smallest. 1/2, 2/3, 2/5, 2/9

Answer. 2/9



Level 2. Routine procedure

Question .

A sales promotion in a store advertises a jacket regularly priced at $125. but now 

selling for 20% off the regular price. You have $100 in 

your pocket (or $100 left in your charge account). Do you have enough money to buy 

the jacket?

Answer .

Yes the money is enough because

New selling price is. 125 x (80/100) =$100.






Level 3. Complex procedures

Question. Solve (3 – 2i)(9 + 4i)


Answer.(3-2i)(9+4i)= 27 +12i–18i – 8i2, which is the same as 27 – 6i – 8(–1), or 35 – 6i.



Level 4 Problem Solving

Question. Karen is thinking of a number. If you double it, and subtract 7, you obtain 11. What is Karen’s number?


Answer

1. We start with 11 and work backwards.

2. The opposite of subtraction is addition. We will add 7 to 11. We are now at 18.

3.The opposite of doubling something is dividing by 2: 18/2 = 9

4.This should be our answer. Looking back:

9 x 2 -7 = 18 -7 = 11

5. We have the right answer.